OpenStudy (anonymous):

Traffic signs are regulated by the Manual on Uniform Traffic Control Devices (MUTCD). The perimeter of some traffic must be 80 inches. Also, its length is 10 inches longer than its width. Find the dimensions of this sign The width of the sign is____inch The length of the sign is ____inch

OpenStudy (anonymous):

let "L" be the length and "W" the width. You need to translate the problem to this notation "The perimeter of some traffic must be 80 inches"--->L+W+L+W=2L+2W=80 which means L+W=40 "its length is 10 inches longer than its width"---->L=W+10--->L-W=10 Now you have an euqtions system: L+W=40 L-W=10 Add both equations and get L+L+W+(-W)=40+10--->2L=50--->L=25 and W=40-L=15 Check it! Perimeter = 2(L+W)=2(25+15)=2·40=80 Length-Width=L-W=25-15=10 Is it clear?

OpenStudy (anonymous):

yes thank you again

OpenStudy (anonymous):

Rule of the thumb: Always read carefully the problem statement, define the notation you are going to use and translate the statement into the mathematical notation

OpenStudy (anonymous):

now with the next one Traffic signs are regulated by the Manual on Uniform Traffic Control Devices (MUTCD). The perimeter of some traffic must be 116 inches. Also, its length is 8 inches longer than its width. Find the dimensions of this sign The width of the sign is____inch The length of the sign is ____inch

OpenStudy (anonymous):

You tell me

OpenStudy (anonymous):

Gimme the equations system

OpenStudy (anonymous):

figuring it out I hate math but i have to take it I got a new question because i put it in wrong it entered it backwards

OpenStudy (anonymous):

2x + 2y = 116 sub for y ... (x + 12) 2x + 2 (x + 12) = 116 4x + 24 = 116 4x = 92

OpenStudy (anonymous):

why sub for y ...x+12 ? The problem states x+8 (i guess x is width and y stands for length)